OpenStudy (superhelp101):

4. What is the possible discriminant of the graph? a parabola opening up with vertex near negative 1.2 comma negative 9.1. a y intercept of negative 6 -11 zero 25 73 5. What is the possible discriminant of the graph? a parabola opening up x intercepts near negative 7 and negative 3 -13 zero 15 16

OpenStudy (superhelp101):

@phi @aaronq

OpenStudy (superhelp101):

@Zarkon @dan815

OpenStudy (superhelp101):

@radar @iambatman

OpenStudy (superhelp101):

@e.mccormick

OpenStudy (superhelp101):

@shamil98 @radar plz

OpenStudy (superhelp101):

@zzr0ck3r

OpenStudy (superhelp101):

@Kainui pretty plz i need help

OpenStudy (imstuck):

For 5 i got 16...still working on 4. I know 5 is correct.

OpenStudy (superhelp101):

Really are u sure about 5

OpenStudy (superhelp101):

how did u get 16?

OpenStudy (imstuck):

I am sure about 5. The value of the discriminant comes out to 16. Use the x intercepts to find the quadratic that these two intercepts create, like this:

OpenStudy (imstuck):

\[(x+3)(x+7)=x ^{2}+10x+21\]The discriminant part of that is \[(10)^{2}-4(1)(21)=16\]The math doesn't lie.

OpenStudy (mathmate):

@superhelp101 You need to post the graph. Depending on the value of the leading coefficient, the value of the discriminant varies. If a is the leading coefficient, the discriminant is \(16a^2\).

OpenStudy (imstuck):

For #4, here's what I did: I figured the quadratic from the x^2 part of the equation, which was (x+ 1.2)^2. FOILing that out gave an equation of \[x ^{2}+2.4x+1.44\]The discriminant part of that is:\[(2.4)^{2}-4(1)(1.44)=0\]

OpenStudy (imstuck):

Can't read the x intercepts on graph 1. Can you tell me what they are?

OpenStudy (superhelp101):

(-3.306, 0) and (.806,0)

OpenStudy (mathmate):

For question 2, f(x)=a(x+3)(x+7) and f(-5)=-4, so a = 1, yes, @IMStuck has the right answer of discriminant = 16(1^2)=16.

OpenStudy (imstuck):

Ok, then, redo on the #4...

OpenStudy (superhelp101):

Alright

OpenStudy (imstuck):

The quadratic equation for #4 is \[x ^{2}+2.5x-2.6646\]and the discriminant part for that is:

OpenStudy (imstuck):

\[(2.5)^{2}-4(1)(-2.6646)=16.89\]That one I'm not seeing in your answers. Maybe @mathmate could lend his eyes to this?

OpenStudy (superhelp101):

oh. Could you help me @mathmate

OpenStudy (mathmate):

Try f(x)=a(x-0.806)(x+3.306) and use f(-1.25)=-9.1 or f(0)=-6 to find a = 2.2517 You should get discriminant equal to 78.73 or 85.76 (instead of 73). I believe we discussed this one before.